OpenStudy (anonymous):

I need Help!!! i will give you a medal if you help me! Compare the functions shown below: f(x) = (x + 3)2 - 2 g(x) linear graph with y intercept of negative 1 over 2 and x intercept of 3 h(x) x y -3 2 -2 7 -1 14 0 23 1 34 2 47 3 62 What is the correct order of the functions from least to greatest according to the average rate of change on the interval from x = -1 to x = 3?

OpenStudy (anonymous):

these are the choices f(x), g(x), h(x) g(x), f(x), h(x) h(x), g(x), f(x) g(x), h(x), f(x) i think its the second one but im not sure it could be the third one.

OpenStudy (anonymous):

@ilikemath50 can you help me please!!!!!

OpenStudy (anonymous):

I will need to see the actual graph to answer that

OpenStudy (anonymous):

@ilikemath50 ok hold up

OpenStudy (anonymous):

is this FLVS?

OpenStudy (anonymous):

I really think its b but i would like to know what you get.

OpenStudy (anonymous):

this is a hard one.. @precal @waterineyes @thomaster @Abhisar @abb0t @SolomonZelman

OpenStudy (anonymous):

I don't think I can help with this, sorry :(

OpenStudy (anonymous):

@ilikemath50 its ok thanks though.

OpenStudy (anonymous):

@aum do you think b is correct?

OpenStudy (aum):

I can't just give out answers but I can help finding the answer. Average rate of change of f(x) on the interval from x = -1 to x = 3 is: \( \Large \frac{f(3) - f(-1)}{(3-(-1))} = \frac{f(3) - f(-1)}{4} \)

OpenStudy (anonymous):

ok, so whats next?

OpenStudy (anonymous):

@aum

OpenStudy (anonymous):

so b wasn't right?

OpenStudy (aum):

Find the equation of h(x). It is a straight line and you are given the x and y intercepts. \(\Large \frac{x}{a} + \frac{y}{b} = 1 \) is the equation of a straight line with x-intercept of 'a' and a y-intercept of 'b'.

OpenStudy (aum):

I haven't worked out the problem. There are several steps involved and we have to go through each step before we can arrive at the answer.

OpenStudy (anonymous):

ohh ok. well h(x) average rate of change is 6 right?

OpenStudy (anonymous):

@aum

OpenStudy (aum):

For g(x) average rate of change is:1/6

OpenStudy (anonymous):

ok so what do we do next? @aum

OpenStudy (aum):

Did you find average rate of change for f(x) first?

OpenStudy (anonymous):

no i didn't. @aum

OpenStudy (anonymous):

how do i find it?

OpenStudy (aum):

That should be done first before we can proceed to the next step.

OpenStudy (aum):

I have given the formula in my very first reply.

OpenStudy (anonymous):

so is it f(4)/4 which equals 1? @aum

OpenStudy (aum):

f(x) = (x + 3)^2 - 2 Find f(3) by putting x = 3 on the right hand side. Find f(-1) by putting x = -1 on the right hand side. Then plug the numbers into the formula in my first reply.

OpenStudy (anonymous):

ok hold on.

OpenStudy (anonymous):

f(3)= 34 and f(-1) = 0 right?

OpenStudy (anonymous):

f(-1) =2

OpenStudy (aum):

correct. f(3) = 34; f(-1) = 2. plug it into the formula

OpenStudy (anonymous):

yay! lol ok so is it 26?

OpenStudy (aum):

(34 - 2) / 4 = 32 / 4 = 8.

OpenStudy (anonymous):

i had just got that. sorry i had did it wrong.

OpenStudy (aum):

For the second function, simply find the slope of the straight line. That will be the average rate of change. In this case, slope = y-intercept / x-intercept = ?

OpenStudy (anonymous):

ok what would be the y-intercept though?

OpenStudy (aum):

For g(x) the problem says: "linear graph with y intercept of negative 1 over 2 and x intercept of 3"

OpenStudy (anonymous):

the x would be 3 right?

OpenStudy (aum):

y-intercept = -1/2 x-intercept = 3 y-intercept / x-intercept = ?

OpenStudy (anonymous):

ok hold on

OpenStudy (anonymous):

@aum i got 0.16

OpenStudy (anonymous):

-0.16

OpenStudy (aum):

Yes, -0.16. For the last function, read the numbers off the table. Find h(3). Find h(-1) Plug it into the formula in my first reply (except f(x) now refers to h(x) in the formula)

OpenStudy (anonymous):

is it 12 @aum

OpenStudy (aum):

Correct! So we have the following average rate of change: f(x): 8 g(x): -0.16 h(x): 12 Arrange from lowest to highest.

OpenStudy (anonymous):

whoaaaaa! lol ok so its g(x), f(x), h(x) @aum

OpenStudy (aum):

Correct!

OpenStudy (aum):

And I did not find out the answer any sooner than you did.

OpenStudy (anonymous):

so it was b! lol Thank You so much!!!!!

OpenStudy (aum):

You are welcome.

OpenStudy (anonymous):

aww thanks for the medal, you are good at what you do!!!

OpenStudy (anonymous):

@aum

OpenStudy (aum):

Oh, thanks. Glad to be able to help.